OpenStudy (anonymous):

Find the dimensions and maximum area of a rectangle whose perimeter is 24 inches.

OpenStudy (anonymous):

Okay so rectangle has two 'lengths' and two 'widths' [ie four sides make up the rectangle] the length cannot be the same as the width other wise it would be a square! So you know 2*length + 2*width = 24 [the perimeter is the measurement of the outer edge] so if we call length 'y' and width 'x' then 2*y + 2*x =24 then you can put in different numbers to see what fits [e.g for a rectangle of perimeter 20. length =6, width = 4 as the perimeter would equal 6+6+4+4=20] The area is then calculated by length*width So in this example for perimeter =20 length =6 width =4 Area=length*width = 6*4 =24 Then apply this to your own example Hope this helps :)